OpenStudy (anonymous):

solve using elimination method. 5x+6y=3 and 10x+12y=6. I have no variables that are the same and i dont get this equation

OpenStudy (bahrom7893):

x and y are the same.. eliminate either one: 5x + 6y = 3 10x + 12y = 6

OpenStudy (bahrom7893):

now solve for x in the first equation: 5x = 3 - 6y x = (3-6y)/5

OpenStudy (radar):

That is an identity.

OpenStudy (anonymous):

I thought they had to be the same variable

OpenStudy (bahrom7893):

plug that into the 2nd equation: 10 * [ (3-6y)/5 ] + 12y = 6 2 * (3-6y) + 12y = 6

OpenStudy (bahrom7893):

divide everything by 2: 3-6y + 6y = 3

OpenStudy (anonymous):

can i multiply the first equation by 2 and the secon by 1

OpenStudy (bahrom7893):

3 = 3, I'm getting infinitely many solutions..

OpenStudy (bahrom7893):

radar check my work.. im goin crazy here..

OpenStudy (radar):

It is an identity, there is no specific solution,

OpenStudy (anonymous):

yea i was getting different answers

OpenStudy (radar):

The two equations do not provide info for a specific solution, as the two equations are identical.

OpenStudy (anonymous):

so im guessing there is infiniely solutions?

OpenStudy (anonymous):

thanks soo much. i will def. be back for more help.